# Bayesian Model Parameter Learning in Linear Inverse Problems: Application in EEG Focal Source Imaging

### v1

@@ 7 Conclusion and Future Work
As an application, we studied the EEG source imaging problem in which the (forward) leadfield model contained a non-linear unknown model parameter, the skull conductivity. The utilization of the BAE approach clearly improved the source localization accuracy of the imaging and the GP provided superior estimates for the skull conductivity when compared to the other tested approaches.
In the future, the BAE approach with GPs will be tested with experimental EEG data, and the estimated skull conductivities will be compared to results from other (experimental) skull conductivity calibration techniques. In EEG imaging, the BAE approach could be tested with more complicated priors, such as $\ell_{1}$ and group-lasso priors, and in scenarios where several tissue conductivities are recovered simultaneously with the source configuration. Finally, as the proposed developments are not limited to the EEG source imaging, other inverse problems such as kernel estimation in blind deconvolution problems could be studied. —————————————————————————————————
@@ Acknowledgements
Appendix A1: Predicting $\sigma$ under Linear Approximation
@@
In this section, we try to explain why the estimates performed well only under these circumstances. First, we try to derive an explicit expression for the posterior $\pi(\sigma|\alpha_{l})$ using the sampling distributions and a linear approximation of the approximation error.
@@
$\sigma_{\mathrm{MAP}}=\sigma_{*}+k_{l}^{-1}(\alpha_{l}+c)\frac{1}{\sqrt{2}\gamma},$
@@
Hence, this modeling demonstrates good predictive capabilities only when the source amplitude is close to the mode (as was shown with simulations in Section 6.1). Therefore, in general, more advanced methods that take into account the estimated source amplitude work better (as was also shown in Section 6.1).
Appendix A2: Choice of the GP Function

### v2

@@ 6 Conclusion and Future Work
As an application, we studied the EEG source imaging problem in which the (forward) leadfield model contained a non-linear unknown model parameter, the skull conductivity. The utilization of the BAE approach clearly improved the source localization accuracy of the imaging. The proposed CG+Iter. and GP provided superior estimates for the skull conductivity when compared to the CG estimates. While the GP and CG+Iter. yielded comparable results across different source amplitudes, the GP method offers a practical advantage due to its computational efficiency and robustness against local minima, making it a preferable choice in most scenarios.
In the future, the BAE approach with GPs will be tested with experimental EEG data, and the estimated skull conductivities will be compared to results from other (experimental) skull conductivity calibration techniques. In EEG imaging, the BAE approach could be tested with more complicated priors, such as $\ell_{1}$ and group-lasso priors, and in scenarios where several tissue conductivities are recovered simultaneously with the source configuration. However, in these new formulations, the complexity of the modeling error (e.g., existence of higher-order frequencies and thus the requirement to use more eigenvectors to describe it) should be investigated. Finally, as the proposed developments are not limited to the EEG source imaging, other inverse problems such as kernel estimation in blind deconvolution problems could be studied.
@@ Acknowledgements
ORCID iDs
Alexandra Koulouri https://orcid.org/$0$-0001-7605-3844
Ville Rimpiläinen https://orcid.org/$0$-0001-8679-7898
Appendix A1: Expressing Approximation Error via Basis Functions
Our aim is to express the approximation error term with the help of a small set of basis functions. We begin with the covariance of the approximation error, defined as
$\Gamma_{\varepsilon}=\mathbb{E}[(\varepsilon-\varepsilon_{*})(\varepsilon-\varepsilon_{*})^{\mathrm{T}}],$ (38)
where $\varepsilon_{*}$ is the mean of the approximation error. Performing eigenvalue decomposition of $\Gamma_{\varepsilon}$ yields
$\Gamma_{\varepsilon}=\sum_{k=1}^{m}\lambda_{k}w_{k}w_{k}^{\mathrm{T}}=W\Lambda W^{\mathrm{T}},$ (39)
where $\lambda_{k}\geq 0$ are the eigenvalues, and $w_{k}\in\mathbb{R}^{m}$ are the corresponding orthonormal eigenvectors (principal components) [43, 42].
This decomposition allows us to express the centered approximation error $\varepsilon-\varepsilon_{*}$ as a linear combination of the eigenvectors:
$\varepsilon-\varepsilon_{*}\in\mathrm{span}\{w_{1},\ldots,w_{m}\}.$ (40)
In practice, we can write the error as
where $\varepsilon^{\prime}=\sum_{k=1}^{p}\alpha_{k}w_{k}$ captures the dominant part of the variation, and $\varepsilon^{\prime\prime}=\sum_{j=p+1}^{m}\beta_{j}w_{j}$ is the residual component. Both components are zero-mean:
$\mathbb{E}[\varepsilon^{\prime}]=\mathbb{E}[\varepsilon^{\prime\prime}]=0.$ (42)
The coefficients $\alpha_{k}$ and $\beta_{j}$ are obtained through inner products:
$\alpha_{k}=\langle\varepsilon-\varepsilon_{*},w_{k}\rangle,\quad\beta_{j}=\langle\varepsilon-\varepsilon_{*},w_{j}\rangle.$ (43)
Thus, the approximation error can be compactly represented using a reduced basis $W^{p}=[w_{1},\ldots,w_{p}]\in\mathbb{R}^{m\times p}$ as
$\varepsilon^{\prime}=W^{p}\alpha,\quad\alpha=(\alpha_{1},\ldots,\alpha_{p})^{\mathrm{T}}\in\mathbb{R}^{p}.$ (44)
The residual term can be written as
$\varepsilon^{\prime\prime}=Q\beta=QQ^{\mathrm{T}}(\varepsilon-\varepsilon_{*}),$ (45)
where $Q=[w_{p+1},\ldots,w_{m}]\in\mathbb{R}^{m\times(m-p)}$ spans the orthogonal complement. The variances of the principal coefficients are
$\mathrm{var}(\alpha_{k})=\lambda_{k},\quad\text{for }k=1,\ldots,p.$ (46)
Appendix A2: Predicting Skull Conductivity $\sigma$ under Linear Approximations
@@
In this section, we explain why the estimates performed well only under these circumstances. First, we try to derive an explicit expression for the posterior $\pi(\sigma|\alpha_{l})$ using sampling distributions and a linear approximation for the approximation error.
@@
$\hat{\sigma}_{\mathrm{MAP}}=\sigma_{*}+k_{l}^{-1}(\alpha_{l}+c)\frac{1}{\sqrt{2}\gamma},$
@@
Therefore, the CG modeling demonstrates good predictive capabilities only when the source amplitude is close to the mode (as was shown with simulations in Section 5.2.1), and more advanced methods that take into account the estimated source amplitude work better in general cases (as was also shown in Section 5.2.1).
Appendix A3: Choice of the GP Function for Skull Conductivity Prediction in the EEG Problem
@@
Appendix A4: Pipeline for Model Parameter Estimation Using the BAE Approach
In this section, we describe an example pipeline to set up the BAE approach with combined estimation of primary unknowns and model parameters. We wish to emphasize that the implementation details are problem-specific.
1. Choose an approximate (discrete) observation model with model parameter $\sigma_{0}$. The corresponding model for the inversion is then of the form: $v=A(\sigma_{0})x+W^{p}\alpha+\nu\in\mathbb{R}^{m}$, where $\nu=e+\varepsilon_{*}+\varepsilon^{\prime\prime}$ includes mean of the (additive) measurement noise $e_{*}$, mean of the approximation error term, and the remainder of the approximation error term. Here, the approximation error has been decomposed as $\varepsilon=W^{p}\alpha+\varepsilon_{*}+\varepsilon^{\prime\prime}$. The BAE approach can be used to take into account unknown/uncertain model parameters and numerical (discretization) errors due to model reduction mappings, for example.
2. Measurement noise $e$: Usually in calibration and controlled/repeated experiments the level of noise can be low. If possible, the level of measurement noise should be compared to the approximation error term. For example, one could expect the BAE approach to improve the results when $\|e_{*}\|^{2}+\mathrm{trace}\,\Gamma_{e}<\|\varepsilon_{*}\|^{2}+\mathrm{trace}\,\Gamma_{\varepsilon}$ [42]. In some cases, when $e_{*}(k)^{2}+\Gamma_{e}(k,k)<\varepsilon_{*}(k)^{2}+\Gamma_{\varepsilon_{*}}(k,k)$ holds even for few values of $k$, the BAE approach might still improve the results [42]. Note that here additive measurement noise has been assumed.
3. Model-based learning:
  • Estimate the parameters of the model approximation error, $\varepsilon=A(\sigma)x-A(\sigma_{0})x$, as follows:
    1. 1.
      Draw $S+K$ samples: $x^{(s)}\sim\pi(x)$ and $\sigma^{(s)}\sim\pi(\sigma)$
    2. 2.
      Estimate sample models: $\{A(\sigma^{(1)}),\ldots,A(\sigma^{(S+K)})\}$
    3. 3.
      Evaluate approximation error samples: $\varepsilon^{(s)}=A(\sigma^{(s)}){x}^{(s)}-A_{0}x^{(s)}$
    4. 4.
      Estimate the mean $\varepsilon_{*}$ and covariance matrix $\Gamma_{\varepsilon}$ using the subset $\{x^{(s)},\sigma^{s},\varepsilon^{(s)}\}_{s=1}^{S}$ as
      $\varepsilon_{*}=\frac{\sum_{s=1}^{S}{\varepsilon}_{i}^{(s)}}{S}\,\,\mbox{and}\,\,\Gamma_{\varepsilon}=\frac{1}{S-1}\sum_{s=1}^{S}(\varepsilon^{(s)}-\varepsilon_{*})(\varepsilon^{(s)}-\varepsilon_{*})^{\mathrm{T}}.$
    5. 5.
      Estimate basis functions $W$ with the help of, for example, eigenvalue decomposition of $\Gamma_{\varepsilon}=\Sigma_{k=1}^{m}\lambda_{k}w_{k}w_{k}^{\mathrm{T}}=W\Lambda W^{\mathrm{T}}$, where $\lambda_{1}>\ldots>\lambda_{m}$ and $W=[w_{1},\ldots,w_{m}]$.
    6. 6.
      Select the number of basis functions $W^{p}=[w_{1},\ldots,w_{p}]$ needed for model parameter inference based on the complexity of the model approximation error. The complexity can be reflected in the distribution of the eigenvalues. The choice of $p$ is problem specific, and the analysis of the (semi)analytical expression of the error covariance can be helpful, as presented in this paper. For example, one could check whether $\varepsilon\propto x\sigma$ and whether any high-order terms are needed to describe the variations caused by the model parameter.
    7. 7.
      Evaluate the samples of $\alpha^{(s)}=(W^{p})^{\mathrm{T}}(\varepsilon^{(s)}-\varepsilon_{*})$ with mean $\alpha_{*}=0$ and covariance $\Gamma_{\alpha}=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{p})$.
    8. 8.
      The remaining error is $\varepsilon^{\prime\prime(s)}=QQ^{\mathrm{T}}(\varepsilon^{(s)}-\varepsilon_{*})$, where $Q=[w_{p+1},\ldots,w_{m}]$ with mean $\mathbb{E}[\varepsilon^{\prime\prime}]=0$ and covariance $\Gamma_{\varepsilon^{\prime\prime}}$.
4. Design the optimization and, if needed, choose appropriate prior for the primary unknown based on the properties of the problem:
$(\hat{x},\hat{\alpha})\leftarrow\min_{x,\alpha}\left\{\|\frac{1}{2}L_{\nu}(A_{0}x+W^{p}\alpha+\nu_{*}-v)\|_{2}^{2}-\ln\pi(x,\alpha)\right\},$
where $\nu_{*}=\varepsilon_{*}+e_{*}$ and $\Gamma_{\nu}=\Gamma_{e}+\Gamma_{\varepsilon^{\prime\prime}}$. For $\alpha$, one can often use a zero-mean Gaussian smoothness prior.
5. Model parameter $\sigma$ inference: This step is carried out with the help of the solved $\hat{\alpha}$ from the previous step. In some cases, one may also need the values of $\hat{x}$ for the inference of $\sigma$.
  • For the following cases, it is often sufficient to utilize the conditional Gaussian (CG) approximation, $\hat{\sigma}_{\mathrm{CG}}=\sigma_{*}+\Gamma_{\sigma\alpha}\Gamma_{\alpha}^{-1}\hat{\alpha}$ where $\Gamma_{\sigma\alpha}$ is estimated from samples $\sigma^{(s)}$ and $\alpha^{(s)}$. Feasible model parameter estimates can be expected when
    1. 1.
      Operating close to the linearization point, or in other words, when one can assume that the true value of $\sigma$ is close to the chosen approximate $\sigma_{0}$ that is used in the standard model and the relationship between $\sigma$ and $\alpha$ is linear (or almost linear).
    2. 2.
      The interaction effect of the primary unknown $x$ on $\sigma$ is only modest.
  • In other cases, the model parameter $\sigma$ should be inferred using both $(\hat{\alpha},\hat{x})$. In the following options, the mapping $(\alpha,x)\rightarrow\sigma$ should be smooth and invertible.
    1. 1.
      Option 1: In some cases, such as in this paper, a nonlinear iterative algorithm can be incorporated and initialized with the CG approximation. Iterative algorithms can be time-consuming and computationally effortful. They do not usually require any training data, but then again, they do not benefit from the support of the training data either.
    2. 2.
      Option 2: A Gaussian Process can be constructed based on the pre-computed samples $\{x^{(s)},\sigma^{(s)},\alpha^{(s)}\}_{s=S+1}^{K+S}$. The mean and covariance functions can be chosen with the help of the underlying physics of the problems and/or the training data. To perform well, the training data should cover a broad range of $\sigma$ values. GPs can be simple and fast to compute, but require an additional training step.